# Post-detection inference for sequential changepoint localization

### v1

@@ Appendix C A heuristic method for composite pre- or post-change
C.1 Known pre-change and unknown post-change distributions
@@
$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
@@
15      if $M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
@@ 16            $\mathcal{C}=\mathcal{C}\cup\{t\}$
C.2 Composite pre- and post-change distributions
@@
$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathbbm{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
@@ 17            $\mathcal{C}=\mathcal{C}\cup\{t\}$
C.3 Experimental Results
C.3.1 Known pre-change and composite post-change (Setting II)
We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$. For detection, we employ a weighted CUSUM detector (3.16) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 5 to construct confidence sets. Table C.1 summarizes the results, including the average size and the conditional and unconditional coverage across $100$ independent runs. For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
Table C.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.16) with threshold $A=1000$.
@@ C.3.1 Known pre-change and composite post-change (Setting II)
500 $N(0,1)$$\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.95 0.86 14.06
500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.85 20.83
C.3.2 Composite pre- and post-change (Setting III)
Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$. We use a weighted CUSUM detector with RIPr defined in (4.17) for change detection. The thresholds are set at $A=1000$ for both $T=100$ and $T=500$. We considered the same discrete weight distribution as above, taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=0.8,d=0.2$ and for Poisson distribution $\theta=1.8,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 6 to construct confidence sets. We report the average size and the conditional and unconditional coverage across $100$ independent runs in Table C.2. The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
Table C.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 6 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
@@ C.3.2 Composite pre- and post-change (Setting III)
Appendix D Another heuristic method for composite pre- or post-change
@@
For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section C.3.1 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table D.1. Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section C.3.2 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table D.1. From Table D.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
Table D.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.16) with threshold $A=1000$.
@@
500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.95 0.86 27.17
@@
Appendix E Summary of comparisons of heuristic and non-heuristic methods
We first compare the execution times of the three types of algorithms and present the results in Table E.1. Algorithm 3 and Algorithm 4 are collectively referred to as Proposed Method, which comes with theoretical guarantees but requires more computational time compared to the heuristic approaches. The first heuristic approach, described in Algorithm 5 and Algorithm 6, is referred to as Heuristic 1. Finally, the heuristic introduced in this section (which is Algorithm 2, with $F_{0}$ and $F_{1}$ being the “closest” pair of distributions in the composite pre- and post-change models) is labeled as Heuristic 2.
Table E.1: Average (of $10$ independent runs) execution times (in seconds) taken by different methods, with $N=B=100$, $\alpha=0.05$, and for the proposed method, $\beta=\gamma=0.025$.

### v2

@@ Appendix D A heuristic method for composite pre- or post-change
D.1 Known pre-change and unknown post-change distributions
@@
$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
@@
15      if $M_{t}\leq\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B})$ then
@@ 16            $\mathcal{C}=\mathcal{C}\cup\{t\}$
D.2 Composite pre- and post-change distributions
@@
$M_{t}>\operatorname{Quantile}(1-\alpha\times\sum_{i=1}^{N}\mathds{I}(\tau_{i}\geq t)/N;M_{t},M_{t}^{1},\cdots,M_{t}^{B}).$
@@ 17            $\mathcal{C}=\mathcal{C}\cup\{t\}$
D.3 Experimental Results
D.3.1 Known pre-change and composite post-change (Setting II)
We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$. For detection, we employ a weighted CUSUM detector (3.15) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 5 to construct confidence sets. Table D.1 summarizes the results, including the average size and the conditional and unconditional coverage across $100$ independent runs. For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
Table D.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.15) with threshold $A=1000$.
@@ D.3.1 Known pre-change and composite post-change (Setting II)
500 $N(0,1)$$\{N(\mu,1):\mu>0\}$ $N(1,1)$ 0.95 0.91 14.06
500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.89 20.83
D.3.2 Composite pre- and post-change (Setting III)
Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$. We use a weighted CUSUM detector with RIPr defined in (4.17) for change detection. The thresholds are set at $A=1000$ for both $T=100$ and $T=500$. We considered the same discrete weight distribution as above, taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=0.8,d=0.2$ and for Poisson distribution $\theta=1.8,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 6 to construct confidence sets. We report the average size and the conditional and unconditional coverage across $100$ independent runs in Table D.2. The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
@@ D.3.2 Composite pre- and post-change (Setting III)
Appendix E Another heuristic method for composite pre- or post-change
@@
For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section D.3.1 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table E.1. Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section D.3.2 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table E.1. From Table E.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
Table E.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.15) with threshold $A=1000$.
@@
500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.95 0.90 27.17
Table E.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
@@
Summary of comparisons of heuristic and non-heuristic methods
We first compare the execution times of the three types of algorithms and present the results in Table E.3 for the Gaussian mean shift setting. Algorithm 3 and Algorithm 4 are collectively referred to as Proposed Method, which comes with theoretical guarantees but requires more computational time compared to the heuristic approaches. The first heuristic approach, described in Algorithm 5 and Algorithm 6, is referred to as Heuristic 1. Finally, the heuristic introduced in this section (which is Algorithm 2, with $F_{0}$ and $F_{1}$ being the “closest” pair of distributions in the composite pre- and post-change models) is labeled as Heuristic 2.
Table E.3: Average (of $10$ independent runs) execution times (in seconds) taken by different methods, with $N=B=100$, $\alpha=0.05$, and for the proposed method, $\beta=\gamma=0.05$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(1,1)$ respectively.